# Gradient Descent on Logistic Regression with Non-Separable Data and Large Step Sizes

### v1

@@ A.2 Cycle construction below the critical step size
Since $\ell^{\prime\prime}$ is decreasing on $[0,\infty)$, so is ${\mathcal{L}}^{\prime\prime}$. By continuity, there must exist $w>0$ such that $T^{\prime}(w)<0$, as $\gamma>1$. As $T(0)=0$ and $T^{\prime}(0)=1-\gamma<0$, $T$ is decreasing in the neighbourhood of 00 to the right. Thus there exists $\tilde{w}>0$ depending on the value of $\gamma\in(1,2]$ sufficiently close to 00 such that $T(\tilde{w})<0$.
@@
$\frac{\partial G}{\partial w}(c,\tilde{w})=1-\eta_{0}{\mathcal{L}}_{0}^{\prime\prime}(\tilde{w})\neq 0.$
Define $\phi:\mathbb{R}^{2}\to\mathbb{R}$ to be $\phi(c,w)\coloneqq{\mathcal{L}}_{c}(w)$. Clearly, $\phi$ is twice-continuously differentiable in both both variables as $\ell$ is twice-continuously differentiable. Assumption 1 on $\ell$ lets us invoke Theorem 6 to have that $w^{*}_{c}$ is continuous and differentiable as a function of $c$. Finally, by the chain rule, $T_{c}(w)$ is also continuously differentiable as a function of both $c$ and $w$. Together, we have that $G$ is continuously differentiable in a neighbourhood of the point $c=0$ and $w=\tilde{w}$. By the implicit function theorem (Theorem 4), there exists a function $\omega:I\to J$ where $I$ is an open interval about $c=0$ and $J$ is an open interval about $w=\tilde{w}$ such that $G(c,\omega(c))=0$ for all $c\in I$. It suffices to pick a nonzero and rational $c_{0}\in I$ such that $\bar{w}=\omega(c_{0})>0$ and $G(c_{0},\bar{w})=T_{c_{0}}(\bar{w})-T(\tilde{w})=0$.
@@
which is continuous in $w$ and $\epsilon$ for all $\epsilon\neq 0$ and $w\neq 0$. At the point $\epsilon=0$ and $w\neq 0$, observe that
@@
as long as $w\neq 0$, and so $\frac{\partial^{2}\rho}{\partial\epsilon\partial w}$ is continuous at $\epsilon=0$ as long as $w\neq 0$. Finally, consider $w>0$,
@@
We will invoke Theorem 6 to show that $w^{*}_{\epsilon}$ is a continuous function in $\epsilon$. To verify the assumptions, first observe that as a function of $\epsilon\in\mathbb{R}$ and $w\in\mathbb{R}$, the perturbed loss ${\mathcal{L}}_{\epsilon}$ is clearly proper as it never attains $\infty$- ∞ and its (effective) domain is nonempty. It is also continuous in both $\epsilon$ and $w$ by definition of $\ell_{\epsilon}$ and continuity of ${\mathcal{L}}$. By Assumption 1 $\ell$ is strictly convex, and so is ${\mathcal{L}}_{\epsilon}$ as its the sum of a strictly convex function with a convex function $\ell_{\epsilon}$.
It remains to show that the horizon function (Definition 2) ${\mathcal{L}}^{\infty}_{0}(w)>0$ for all $w\neq 0$. Using the formula for a convex ${\mathcal{L}}_{\epsilon}$ from Theorem 5 we have for all $w\neq 0$ and arbitrary $\bar{x}\in\mathbb{R}$,
@@
is continuously differentiable everywhere except at $w=0$ and $\epsilon=0$.
@@ B.1 Cycle construction in two dimensions
Consider the dataset constructed in Lemma 6 where all labels are $y_{i}=1$, and the $x_{i}\in\mathbb{R}^{2}$’s consist of $m_{1}$ copies of $e_{1}$, $n_{1}$ copies of $-e_{1}$, $m_{2}$ copies of $e_{2}$, and $n_{2}$ copies of $-e_{2}$, where $e_{i}$ is the $i$-th coordinate vector. The resulting loss, gradient, and Hessian are given by
@@
By Lemmas 6 and 7, under the step size $\eta$, there exists a GD trajectory where $w_{0},w_{1},\dots,w_{k-2}$ all lie in the positive quadrant below the line $w^{\mathsf{\scriptscriptstyle T}}p=0$, while $w_{k-1}$ crosses this line but remain to the left of $w^{*}$ in the positive quadrant. Furthermore, next point $w_{k}$ will move closer to $w^{*}$ in both directions without crossing back, due to the decoupling of our objective and that $\gamma\leq 1$.
Now let
$p=(w_{2}^{*},\,-w_{1}^{*})^{\mathsf{\scriptscriptstyle T}}\quad\text{and}\quad q=w_{0}-T(w_{k}+\eta\cdot c_{1}\cdot p),$ (39)
@@
Let $c_{2}=1/\eta$, where $\eta$ is the step size that generated the trajectory in the original trajectory described above. Then for all $\epsilon$ such that $c_{1}\cdot\epsilon^{2}$ and $c_{2}\cdot\epsilon^{2}$ are both rational, the perturbed objective
@@
be the corresponding GD map. Let $\tilde{w}_{t+1}=T_{\epsilon}(\tilde{w}_{t})$ be the GD iterates under this perturbed map $T_{\epsilon}$, starting at $\tilde{w}_{0}=w_{0}$. By continuity of $\eta_{\epsilon}$, $\nabla{\mathcal{L}}_{\epsilon}$, and $\nabla^{2}{\mathcal{L}}_{\epsilon}$ (Lemmas 8 and 9), we have
@@
At $w_{k-1}$ however, the first gradient becomes active, and thus we have
$\lim_{\epsilon\to 0}\tilde{w}_{k}=\lim_{\epsilon\to 0}T_{\epsilon}(\tilde{w}_{k-1})$ $=\lim_{\epsilon\to 0}\tilde{w}_{k-1}-\eta_{\epsilon}\nabla{\mathcal{L}}_{\epsilon}(\tilde{w}_{k-1})$
$=w_{k-1}-\eta\nabla{\mathcal{L}}(w_{k-1})+\eta\cdot c_{1}\cdot p+0$
$=w_{k}+\eta\cdot c_{1}\cdot p.$
@@
$(w_{k}+\eta\cdot c_{1}\cdot p)^{\mathsf{\scriptscriptstyle T}}q=(w_{k}+\eta\cdot c_{1}\cdot p)^{\mathsf{\scriptscriptstyle T}}(w_{0}-T(w_{k}+\eta\cdot c_{1}\cdot p))$
@@
$\lim_{\epsilon\to 0}T_{\epsilon}(\tilde{w}_{k})$ $=\lim_{\epsilon\to 0}\tilde{w}_{k}-\eta_{\epsilon}\nabla{\mathcal{L}}_{\epsilon}(\tilde{w}_{k})$
$=w_{k}+\eta\cdot c_{1}\cdot p-\eta\nabla{\mathcal{L}}(w_{k}+\eta\cdot c_{1}\cdot p)-0+\eta\cdot c_{2}\cdot q$
$=T(w_{k}+\eta\cdot c_{1}\cdot p)+\eta\frac{1}{\eta}q$
$=T(w_{k}+\eta\cdot c_{1}\cdot p)+w_{0}-T(w_{k}+\eta\cdot c_{1}\cdot p)$
@@
$G(\epsilon,w)\coloneqq T^{k+1}_{\epsilon}(w)-w=\underbrace{T_{\epsilon}\circ\ldots\circ T_{\epsilon}}_{k+1\text{ times}}(w)-w.$ (44)
@@
$J_{w}G(\epsilon,w)$ $=J_{w}T_{\epsilon}^{k+1}(w)-I$
$=\prod_{t=0}^{k}\left(I-\eta_{\epsilon}\nabla^{2}{\mathcal{L}}_{\epsilon}(T_{\epsilon}(\tilde{w}_{t}))\right)-I.$
@@
$\left|\lambda_{i}\left(J_{w}T_{0}^{k+1}(w_{0})\right)\right|<(1-\gamma)^{k}<1.$
That is, all eigenvalues of the Jacobian of $T^{k+1}$ lies within the unit circle, when $\epsilon=0$ and $w=w_{0}$. This suffices to ensure that
@@
$\mu(\epsilon,w)$ $\coloneqq\left|(J_{w}T_{\epsilon}^{k+1})(w)\right|$ (45)
@@
We know that $\mu(0,w_{0})<1$ from earlier, and by continuity there must exist a neighbourhood $U$ around $\epsilon=0$ such that $\mu(\epsilon,\omega(\epsilon))<1$ for all $\epsilon\in U$, as $\omega$ is continuous in $\epsilon$. Since $U$ and $I$ are both open, there must exist a nonzero $\epsilon_{0}\in I\cap U$ such that
@@
implying that the trajectory on the perturbed objective with $\epsilon=\epsilon_{0}$ is a stable cycle. Stability ensures that if initialize close enough to $\omega(\epsilon_{0})$ we will converge to this cycle, as it is a stable fixed point of the $T^{k+1}$ map. A final note is that we can always pick $\epsilon_{0}$ such that $c_{1}\cdot\epsilon_{0}^{2}$ and $c_{2}\cdot\epsilon_{0}^{2}$ are both rational, so that the perturbed objective 41 indeed corresponds to a valid classification problem. This completes the proof.
@@
where $y_{i}=1$ for all $i$, and there are $m_{1}$ copies of $e_{1}$, $n_{1}$ copies of $-e_{1}$, $m_{2}$ copies of $e_{2}$, and $n_{2}$ copies of $-e_{2}$, with $e_{i}$ being the $i$-th coordinate vector in $2d$. Suppose $\ell$ satisfies Assumption 1, and $w_{1}^{*}>w_{2}^{*}>0$. Then for all $\gamma\in(0,1]$, there exists a combination of $m_{1}$, $n_{1}$, $m_{2}$, and $n_{2}$ such that with a step size of $\eta=\gamma/\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$, there exists a point $w=(w_{1},\,w_{2})^{\mathsf{\scriptscriptstyle T}}$ with
@@
are all continuous everywhere except on the set 49, where $J_{\epsilon}$ denotes the Jacobian with respect to $\epsilon$. Lastly, $\frac{\partial^{2}\rho}{\partial^{2}\epsilon}$ only continuous on $\left\{(\epsilon,w):\,w^{\mathsf{\scriptscriptstyle T}}p>0\right\}$.
@@
Continuity of $w^{*}_{\epsilon}$ follows the same argument as in the one-dimensional case (Lemma 4), where the horizon function ${\mathcal{L}}^{\infty}_{0}(w)>0$ for all $w$ not at the origin follows again from convexity.
Differentiability needs a bit more attention. As before, let
@@
is continuously differentiable everywhere in $\epsilon$, where $\gamma$ is some constant, and ${\mathcal{L}}_{\epsilon}$ is defined as in 41.
@@
Appendix C Derivations for the toy dataset
C.1 Gradient and Hessian
The loss function is given by
with gradient
Applying the chain rule gives us the Hessian
C.2 GD update in probability space
Proposition 1.
$p_{t+1,i}$ $=1-p_{t+1,n}\quad\forall i=1,\dots,n-1.$ (54)
Proof.
$p_{t,i}$ $=\sigma(-y_{i}w_{t}^{\mathsf{\scriptscriptstyle T}}x_{i})$
$=1-p_{t,n}$
and so 3.1 holds. The update for $p_{t,n}$ is
where in the second step we used dataset construction along with 3.1.
Proposition 2.
Proof.
When $n=2$, the one-dimensional map for $p_{n}$ is given by
For a period-$2$ cycle, we have $z_{2}=\varphi(z_{1})$ and $z_{1}=\varphi(z_{2})$, with $z_{1}+z_{2}=1$ (see 55). Therefore, a period-$2$ point $x=z_{1}$ or $x=z_{2}$ must satisfy
which is equivalent to
$\ln\left(\frac{x}{1-x}\right)$ $=\frac{\eta}{2}\left(x-\frac{1}{2}\right).$
Using a change of variable with $u\coloneqq 2x-1$, this is equivalent to
Appendix D Miscellaneous results
D.1 Properties of fhe logistic loss and the squareplus loss
@@
using the fact that $\ell$ is convex with $\ell^{\prime}$ upper bounded by $1$, which gives us the desired limit. ∎
D.2 Technical results in (convex) analysis
Theorem 4 (Implicit function theorem (Krantz and Parks, 2002)).

### v2

@@ A.2 Cycle construction below the critical step size
Since $\ell^{\prime\prime}$ is decreasing on $[0,\infty)$, so is ${\mathcal{L}}^{\prime\prime}$. By continuity, there must exist $w>0$ such that $T^{\prime}(w)<0$, as $\gamma>1$. As $T(0)=0$ and $T^{\prime}(0)=1-\gamma<0$, $T$ is decreasing in the neighborhood of 00 to the right. Thus there exists $\tilde{w}>0$ depending on the value of $\gamma\in(1,2]$ sufficiently close to 00 such that $T(\tilde{w})<0$.
@@
$\frac{\partial G}{\partial w}(0,\tilde{w})=1-\eta_{0}{\mathcal{L}}_{0}^{\prime\prime}(\tilde{w})\neq 0.$
Define $\phi:\mathbb{R}^{2}\to\mathbb{R}$ to be $\phi(c,w)\coloneqq{\mathcal{L}}_{c}(w)$. Clearly, $\phi$ is twice-continuously differentiable in both both variables as $\ell$ is twice-continuously differentiable. Assumption 1 on $\ell$ lets us invoke Theorem 6 to have that $w^{*}_{c}$ is continuous and differentiable as a function of $c$. Finally, by the chain rule, $T_{c}(w)$ is also continuously differentiable as a function of both $c$ and $w$. Together, we have that $G$ is continuously differentiable in a neighborhood of the point $c=0$ and $w=\tilde{w}$. By the implicit function theorem (Theorem 4), there exists a function $\omega:I\to J$ where $I$ is an open interval about $c=0$ and $J$ is an open interval about $w=\tilde{w}$ such that $G(c,\omega(c))=0$ for all $c\in I$. It suffices to pick a nonzero and rational $c_{0}\in I$ such that $\bar{w}=\omega(c_{0})>0$ and $G(c_{0},\bar{w})=T_{c_{0}}(\bar{w})-T(\tilde{w})=0$.
@@
which is continuous in $w$ and $\epsilon$ for all $(\epsilon,w)\neq(0,0)$. At the point $(\epsilon,w)=(0,0)$,
@@
as long as $w\neq 0$, and so $\frac{\partial^{2}\rho}{\partial\epsilon\partial w}$ is continuous at $\epsilon=0$ as long as $w\neq 0$. Finally, for $w>0$,
@@
We will invoke Theorem 6 to show that $w^{*}_{\epsilon}$ is a continuous function in $\epsilon$. To verify the assumptions, first observe that as a function of $\epsilon\in\mathbb{R}$ and $w\in\mathbb{R}$, the perturbed loss ${\mathcal{L}}_{\epsilon}$ is clearly proper as it never attains $\infty$- ∞ and its (effective) domain is nonempty. It is also continuous in both $\epsilon$ and $w$ by definition of $\ell_{\epsilon}$ and continuity of ${\mathcal{L}}$. By Assumption 1, $\ell$ is strictly convex, and so is ${\mathcal{L}}_{\epsilon}$ as it is the sum of a strictly convex function with a convex $\ell_{\epsilon}$.
It remains to show that the horizon function (Definition 1) ${\mathcal{L}}^{\infty}_{0}(w)$ is positive for all $w\neq 0$. Using the formula for a convex ${\mathcal{L}}_{\epsilon}$ from Theorem 5, we have for all $w\neq 0$ and $\bar{x}\in\mathbb{R}$,
@@
is continuously differentiable everywhere except at $(\epsilon,w)=(0,0)$.
@@ B.1 Cycle construction in two dimensions
Consider the dataset constructed in Lemma 6 where all labels are $y_{i}=1$, and the $x_{i}\in\mathbb{R}^{2}$’s consist of $m_{1}$ copies of $e_{1}$, $n_{1}$ copies of $-e_{1}$, $m_{2}$ copies of $e_{2}$, and $n_{2}$ copies of $-e_{2}$, where $e_{i}$ is the $i$th coordinate vector. The resulting loss, gradient, and Hessian are given by
@@
By Lemmas 6 and 7, under the step size $\eta$, there exists a trajectory where $\left\{w_{t}\right\}_{0}^{k-3}$ all lie in the positive quadrant below the line $w^{\mathsf{\scriptscriptstyle T}}p=0$, while $w_{k-2}$ crosses this line but remain to the right of $w^{*}$ in the positive quadrant. Furthermore, next point $w_{k}$ will move closer to $w^{*}$ in both directions without crossing back, due to the decoupling of our objective and that $\gamma\leq 1$.
Now let $w_{1}^{*}$, $w_{2}^{*}$ be the two coordinates of $w^{*}$,
$p=(w_{2}^{*},\,-w_{1}^{*})^{\mathsf{\scriptscriptstyle T}},\quad\text{and}\quad q=w_{0}-T(w_{k-1}+\eta\cdot c_{1}\cdot p),$ (39)
@@
Let $c_{2}=1/\eta$, where $\eta$ is the step size that generated the original trajectory described above. Then for all $\epsilon$ such that $c_{1}\cdot\epsilon^{2}$ and $c_{2}\cdot\epsilon^{2}$ are both rational, the perturbed objective
@@
be the corresponding GD map. Let $\tilde{w}_{t+1}=T_{\epsilon}(\tilde{w}_{t})$ be the GD iterates under this perturbed map $T_{\epsilon}$, starting at $\tilde{w}_{0}=w_{0}$. By continuity of $\eta_{\epsilon}$, $\nabla{\mathcal{L}}_{\epsilon}$, and $\nabla^{2}{\mathcal{L}}_{\epsilon}$ (Lemmas 8 and 9),
@@
$\lim_{\epsilon\to 0}\tilde{w}_{k-1}=\lim_{\epsilon\to 0}T_{\epsilon}(\tilde{w}_{k-2})$ $=\lim_{\epsilon\to 0}\tilde{w}_{k-2}-\eta_{\epsilon}\nabla{\mathcal{L}}_{\epsilon}(\tilde{w}_{k-2})$
$=w_{k-2}-\eta\nabla{\mathcal{L}}(w_{k-2})+\eta\cdot c_{1}\cdot p+0$
$=w_{k-1}+\eta\cdot c_{1}\cdot p.$
@@
$(w_{k-1}+\eta\cdot c_{1}\cdot p)^{\mathsf{\scriptscriptstyle T}}q=(w_{k-1}+\eta\cdot c_{1}\cdot p)^{\mathsf{\scriptscriptstyle T}}(w_{0}-T(w_{k-1}+\eta\cdot c_{1}\cdot p))$
@@
$\lim_{\epsilon\to 0}T_{\epsilon}(\tilde{w}_{k-1})$ $=\lim_{\epsilon\to 0}\tilde{w}_{k-1}-\eta_{\epsilon}\nabla{\mathcal{L}}_{\epsilon}(\tilde{w}_{k-1})$
$=w_{k-1}+\eta\cdot c_{1}\cdot p-\eta\nabla{\mathcal{L}}(w_{k-1}+\eta\cdot c_{1}\cdot p)-0+\eta\cdot c_{2}\cdot q$
$=T(w_{k-1}+\eta\cdot c_{1}\cdot p)+\eta\frac{1}{\eta}q$
$=T(w_{k-1}+\eta\cdot c_{1}\cdot p)+w_{0}-T(w_{k-1}+\eta\cdot c_{1}\cdot p)$
@@
$G(\epsilon,w)\coloneqq T^{k}_{\epsilon}(w)-w=\underbrace{T_{\epsilon}\circ\ldots\circ T_{\epsilon}}_{k\text{ times}}(w)-w.$ (44)
@@
$J_{w}G(\epsilon,w)$ $=J_{w}T_{\epsilon}^{k}(w)-I$
$=\prod_{t=0}^{k-1}\left(I-\eta_{\epsilon}\nabla^{2}{\mathcal{L}}_{\epsilon}(T_{\epsilon}(\tilde{w}_{t}))\right)-I.$
@@
$\left|\lambda_{i}\left(J_{w}T_{0}^{k}(w_{0})\right)\right|<(1-\gamma)^{k}<1.$
That is, all eigenvalues of the Jacobian of $T^{k}$ lies within the unit circle, when $\epsilon=0$ and $w=w_{0}$. This suffices to ensure that
@@
$\mu(\epsilon,w)$ $\coloneqq\left|(J_{w}T_{\epsilon}^{k})(w)\right|$ (45)
@@
We know that $\mu(0,w_{0})<1$ from earlier, and by continuity there must exist a neighborhood $U$ around $\epsilon=0$ such that $\mu(\epsilon,\omega(\epsilon))<1$ for all $\epsilon\in U$, as $\omega$ is continuous in $\epsilon$. Since $U$ and $I$ are both open, there must exist a nonzero $\epsilon_{0}\in I\cap U$ such that
@@
implying that the trajectory on the perturbed objective with $\epsilon=\epsilon_{0}$ is a stable cycle. Stability ensures that if initialize close enough to $\omega(\epsilon_{0})$ we will converge to this cycle, as it is a stable fixed point of the $T^{k}$ map. A final note is that we can always pick $\epsilon_{0}$ such that $c_{1}\cdot\epsilon_{0}^{2}$ and $c_{2}\cdot\epsilon_{0}^{2}$ are both rational, so that the perturbed objective 41 indeed corresponds to a valid classification problem. This completes the proof.
@@
where $y_{i}=1$ for all $i$, and there are $m_{1}$ copies of $e_{1}$, $n_{1}$ copies of $-e_{1}$, $m_{2}$ copies of $e_{2}$, and $n_{2}$ copies of $-e_{2}$, with $e_{i}$ being the $i$th coordinate vector in $2d$. Suppose $\ell$ satisfies Assumption 1, and $w_{1}^{*}>w_{2}^{*}>0$. Then for all $\gamma\in(0,1]$, there exists a combination of $m_{1}$, $n_{1}$, $m_{2}$, and $n_{2}$ such that with a step size of $\eta=\gamma/\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$, there exists a point $w=(w_{1},\,w_{2})^{\mathsf{\scriptscriptstyle T}}$ with
@@
are all continuous everywhere except on the set 49, where $J_{\epsilon}$ denotes the Jacobian with respect to $\epsilon$. Lastly, $\frac{\partial^{2}\rho}{\partial^{2}\epsilon}$ is only continuous on $\left\{(\epsilon,w):\,w^{\mathsf{\scriptscriptstyle T}}p>0\right\}$ (as in Lemma 3). ∎
@@
Continuity of $w^{*}_{\epsilon}$ follows the same argument as in the one-dimensional case (by Lemma 4), where the horizon function ${\mathcal{L}}^{\infty}_{0}(w)>0$ for all $w$ not at the origin follows again from convexity. Differentiability needs a bit more attention. As before, let
@@
is continuously differentiable everywhere in $\epsilon$, where ${\mathcal{L}}_{\epsilon}$ is defined as in 41.
@@
Appendix C Miscellaneous results
C.1 Properties of the logistic loss and the squareplus loss
@@
using the fact that $\ell$ is convex with $\ell^{\prime}$ upper bounded by $1$, which gives us the desired limit.
C.2 Technical results in (convex) analysis
Throughout our proofs, we extensively utilized the most basic form of the implicit function theorem (Krantz and Parks, 2002, Theorem 1.3.1), stated as follows.
Theorem 4 (Implicit function theorem).